Suppose a die is rolled twice and let
B = {first toss is a 1} D = {second toss is a 1}
Find the requested probability. (Enter the probability as a fraction.)
P(B ∪ D)

Answer by ikleyn(52848)

.
Suppose a die is rolled twice and let
B = {first toss is a 1} D = {second toss is a 1}
Find the requested probability. (Enter the probability as a fraction.)
P(B ∪ D)
~~~~~~~~~~~~~~

(B U D) means "B or D".


It is very well known fact that  P(B) = 1/6,  P(D) = 1/6,  P(B ∩ D) = 1/36.


Next, use the formula for the probability of the union of two events

    P(B U D) = P(B) + P(D) - P(B ∩ D) = 1/6 + 1/6 - 1/36 = 6/36 + 6/36 - 1/36 = 11/36.


ANSWER.  P(B U D) = 11/36.

Solved.